Logiweb bibliography of check

Up Help

[0] check

[1] base

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-06.UTC:07:49:23.045248 = MJD-53892.TAI:07:49:56.045248 = LGT-4656296996045248e-6